OpenStudy (anonymous):

Write the standard form of the equation of the line passing the point. (-1, 4), (3, -2). Explanation? Please :)

terenzreignz (terenzreignz):

Okay, two things we need for a line (most of the time) is a slope, and a point... With two points, you can easily get the slope of the line... do you know how?

terenzreignz (terenzreignz):

To refresh you, if you have two points, (a,b) and (c,d), the slope m of the line is given by... \[\Large m = \frac{d-b}{c-a}\]Note that it doesn't matter if you have it as... \[\Large m = \frac{b-d}{a-c}= \frac{-1(d-b)}{-1(c-a)}= \frac{d-b}{c-a}\] Just be careful... and CONSISTENT :)

OpenStudy (anonymous):

Oh my, I'm sure I do, but I can't remember how. This was taught the beginning of the semester, It's a question on my finals. Does it have to do with subtracting the x and y values?

terenzreignz (terenzreignz):

It has everything to do with subtracting x and y values :) Basically, the slope equals the change in y divided by the change in x... denoted \[\Large m= \frac{\Delta y}{\Delta x}\]

terenzreignz (terenzreignz):

That's not yet the answer, though.... that's just the slope... that said, what IS the slope?

OpenStudy (anonymous):

With this equation, it's equal to 1/2. How can standard form most easily be defined as? "Write in standard form. . ."

terenzreignz (terenzreignz):

Nope, it isn't 1/2 at all.... check it again :)

OpenStudy (anonymous):

Oh, I'm dumb. Thanks for the patience by the way. . Ahhhhhhh, 1/2 being the slope?

terenzreignz (terenzreignz):

Yeah... 1/2 is not the slope... (-1 4) and (3 , -2) remember the formula... \[\Large m = \frac{d-b}{c-a}\]

terenzreignz (terenzreignz):

And your (a,b) is (-1 , 4) while (c,d) is (3,-2)

OpenStudy (anonymous):

Right. -2-4 is -6. 3+1=4. . . . -2/3?

terenzreignz (terenzreignz):

backwards.... it's the difference of the Y's over the difference of the X's :P

OpenStudy (anonymous):

LOL, oh gosh. I feel so dumb. Okay, 3/-2. And from there? y=mx+b?

terenzreignz (terenzreignz):

Yes... we'll get there :D so you have the slope, \[\Large m = -\frac32\] right?

OpenStudy (anonymous):

Yes. Got it!

terenzreignz (terenzreignz):

So, you now have a slope m, and a point (two of them, in fact) You pick one of these points (it doesn't really matter which) Have you picked one? Great. Now, let that point be (h,k) The line would be... \[\Large y = \color{red}m(x-\color{green}h)+ \color{blue}k\]

OpenStudy (anonymous):

y=-3/2(x+ -1) +4. Got it.

terenzreignz (terenzreignz):

Be attentive... \[\Large y = m(x\color{red}{-}h) +k\] That's a MINUS sign there...

OpenStudy (anonymous):

Boo! Sorry. Minor changes. . y=-3/2(x+1)+4 Next?

terenzreignz (terenzreignz):

Better :) \[\LARGE y = -\frac32(x+1) + 4\] All that's left is to simplify :)

OpenStudy (anonymous):

Isn't that in standard form?

terenzreignz (terenzreignz):

Not yet... it has to be in the form \[\LARGE \color{red}y = \color{green}m \color{blue}x + \color{orange}b\]

terenzreignz (terenzreignz):

So, distribute that -3/2, and then simplify... that's all there is to it :)

OpenStudy (anonymous):

So. . .y=-4 3/2x - 1 3/2?? What is the answer? I can figure it out from there and work back wards. Lol

terenzreignz (terenzreignz):

You're confusing me... :/ \[\Large y = -\frac32(\color{green}x+\color{green}1)+4\] I just said distribute the -3/2 over the x and the +1, and then simplify.

OpenStudy (anonymous):

That is -3/2x -3/2 + 4?

terenzreignz (terenzreignz):

Okay... good so far :) \[\Large y = -\frac32x -\frac32 + 4\]Now simplify this part... \[\Large y = -\frac32x \color{blue}{-\frac32 + 4}\]

OpenStudy (anonymous):

y=-3/2x -6

terenzreignz (terenzreignz):

Nope... review your addition and subtraction of fractions :)

OpenStudy (anonymous):

What! I was pretty confident. Lol, what is the answer?

terenzreignz (terenzreignz):

Can't tell you that :P Do it over... c'mon, you can do this, it's just one tiny step more :D

OpenStudy (anonymous):

Aww, thanks for the encouragement. Hmm, y=-3/2+4?

terenzreignz (terenzreignz):

What happened to the x? Come on, it has all boiled down to this... \[\LARGE -\frac32 + 4 = \color{red}?\]

OpenStudy (anonymous):

Lol, my bad. . y=-3/2x + 5 1/2?

terenzreignz (terenzreignz):

Nope... note the minus sign on the "3/2" Besides, you shouldn't use mixed-number fractions, use improper fractions...

OpenStudy (anonymous):

LOL, I am so not feeling this. y=-3/2x -6?

terenzreignz (terenzreignz):

Nope...

OpenStudy (anonymous):

OMG. Thanks for your help. Lol, Good Night :)